A recent survey found that 80% of all people over 60 years of age in a large city wear glasses. a random sample of 6 people over 60years of age is selected.

a. what is the probability that exactly 4 people will wear glasses?
b. what is the probability that at most 5 people will wear glasses?
1
Expert's answer
2019-11-12T11:06:16-0500

That is the binomial distribution


P( 4 people from 6 wear glasses ) = C(6, 4) * 0.84 * 0.22 = 0.245

P( at most 5 from 6 wear glasses ) = 1 - P(6 people from 6 wear glasses ) = 1 - 0.86 = 0.737


where C(n, k) is the binomial coefficient.


Answer:


a) 0.245

b) 0.737